Item1 1points ItemSkipped eBook Print References Check my workCheck My Work button is now disabled Item...

Item1

1points

ItemSkipped

eBook

Print

References

Check my workCheck My Work button is now disabled

Item 1

Item 1 1 points Item Skipped

Ten firms compete in a market to sell product X. The total sales of all firms selling the product are $2,000,000. Ranking the firms’ sales from highest to lowest, we find the top four firms’ sales to be $260,000, $220,000, $150,000, and $130,000, respectively. Calculate the four-firm concentration ratio in the market for product X.
Answer #1

ANSWER:

Four firm concentration ratio = (top four firm sales / total sales) * 100 = ((260,000 + 220,000 + 150,000 + 130,000) / 2,000,000) ) * 100 = (760,000 / 2,000,000) * 100 = 0.38 * 100 = 38%

so the four firm concentration ratio is 38%